The speed of light in cubic zirconia is 1.39x10^8 m/s. What is the index of refraction for this substance
1
Expert's answer
2020-05-29T10:03:48-0400

The index of refraction of a zircon

"n=\frac{c}{v}=\frac{3.00\times 10^8}{1.39\times 10^8}=2.16"